OpenStudy (anonymous):

How do I find the derivative of 3^(4x)?

zepdrix (zepdrix):

Hey there :) Here is our exponential derivative as a reminder:\[\large\rm \frac{d}{dx}a^x=a^x(\ln a)\]

zepdrix (zepdrix):

For our problem here, since the exponent is more than simply x, we have one extra step, `chain rule`.

zepdrix (zepdrix):

If you haven't learned your exponential derivative yet, then it's possible your teacher wants you to use `logarithmic differentiation`. Have you seen the previously mentioned rule before? :)

OpenStudy (anonymous):

Therefore the derivative would be 4ln(3)x3^(4x)?

zepdrix (zepdrix):

Hmm looks very close. I'm not sure about that x squished in the middle.\[\large\rm \frac{d}{dx}3^{4x}\quad=\quad 3^{4x}(\ln3)(4x)'\]\[\large\rm \frac{d}{dx}3^{4x}\quad=\quad 3^{4x}(\ln3)(4)\]

OpenStudy (solomonzelman):

Suppose you have, \(\color{#0000ff}{ \displaystyle y=a^x }\) Take the natural-log of both sides. \(\color{#0000ff}{ \displaystyle \ln y=\ln a^x }\) The right side simplifies to. \(\color{#0000ff}{ \displaystyle \ln y=x(\ln a) }\) When you differentiate, remember that: \(\color{}{ [1]\quad d/dx (\ln x)=1/x }\) \(\color{}{ [2]\quad d/dx (\ln f(x))=f'(x)/f(x) }\) (in [2], Chain Rule is applied) (You need a chain rule for y, since it's a function of x, just as f(x) is, and x(ln a), is just x multiplied times some constant ln(a). ) When differentiating both sides, you get: \(\color{#0000ff}{ \displaystyle \frac{y'}{y}=\ln(a) }\) Multiply both sides times y. \(\color{#0000ff}{ \displaystyle y'=y\cdot\ln(a) }\) You know from the beginning that, \(\color{}{ y=a^x }\). Plug that! \(\color{#0000ff}{ \displaystyle y'=a^x\cdot\ln(a) }\) Therefore, you can conclude that: \(\color{#ff0000}{ \displaystyle \frac{d}{dx}(a^x)=a^x\ln(a) }\) ********************************************************* Suppose you had, \(\color{#0000ff}{ \displaystyle y=a^{g(x)} }\) Then, you can apply the same logarithmic differentiation.... \(\color{#0000ff}{ \displaystyle \ln y=g(x)\cdot \ln (a ) \quad \Longrightarrow \quad y'/y=a\cdot g'(x) \quad \Longrightarrow \quad y'=y\cdot a\cdot g'(x) }\) \(\color{#0000ff}{ \displaystyle \Longrightarrow \quad y'=a^{g(x)} \cdot a\cdot g'(x) }\) So, you may conclude: \(\color{#ff0000}{ \displaystyle \frac{d}{dx}(a^{g(x)})=a^{g(x)}g'(x)\ln(a) }\)

OpenStudy (solomonzelman):

I would never actually do these proves myself, this is just differentiation, except that you don't know what the function and the constants are. (I do proves in Physics, because I am bad at it.)